# Two-Photon Absorption in Silicon Using Real Density Matrix Approach

### v1

@@ II.2 Discrete states
$\left(\frac{\partial Y^{(1)}}{\partial t}\right)_{{\rm irrev}}=-\frac{1}{T_{2}}Y.$ (18)
In the discussion of nonlinear effects we also take into account the non-resonant parts of the amplitude $Y$. The excitonic density $Y$ will consists of two parts, $Y_{a},Y_{b}$. Therefore, Eq.(17) generates four equations: a pair for an amplitude $Y_{a}$: one for $Y_{a-}^{(1)}\,\propto\exp(-{i}\omega_{a}t)$, and the second for $Y^{(1)}_{a+}\,\propto\exp({i}\omega_{a}t)$
${i}\hbar\left(i\omega_{a}+\frac{1}{T_{2}}\right)Y^{(1)}_{a+}-H_{eh}Y^{(1)}_{a+}=-{\bf M}{\bf E}_{a}^{*}({\bf R},t),$
${i}\hbar\left(-{i}\omega_{a}+\frac{1}{T_{2}}\right)Y^{(1)}_{a-}-H_{eh}Y^{(1)}_{a-}=-{\bf M}{\bf E}_{a}({\bf R},t),$
and similar equations for $Y_{b\pm}^{(1)}$. In the following, we consider only one component of the vectors ${\bf E}$, ${\bf P}$ and ${\bf M}$.
@@
$Y^{(1)}_{da-}=E_{a}({\bf R},t)\sum_{n}\frac{c_{n}\varphi_{n}({\bf r})}{\hbar(\Omega_{n}-\omega_{a}-{i}/T_{2n})},$
$Y^{(1)}_{da+}=E_{a}^{*}({\bf R},t)\sum_{n}\frac{c_{n}\varphi_{n}({\bf r})}{\hbar(\Omega_{n}+\omega_{a}-{i}/T_{2n})},$ (20)
and similar expressions for $Y_{db\pm}^{(1)}$. The subscript “d” indicates the case of discrete excitonic states. In equations (20), we define
$c_{n00}\equiv c_{n}=\int{d}^{3}rM({\bf r})\varphi_{n00}({\bf r}),$
$\hbar\Omega_{n00}=\hbar\Omega_{n}=E_{g}+E_{n}(\gamma_{a}),$ (21)
$\varphi_{n\ell m}=R_{n\ell}(r)Y_{\ell m}(\theta,\phi),$
where $R_{n\ell}$ are the hydrogen radial functions of an anisotropic Schrödinger equation, $Y_{\ell m}(\theta,\phi)$ are spherical harmonics, $E_{n\ell m}$ the corresponding eigenvalues
$\phantom{nuc}E_{n\ell m}=-\frac{\eta^{2}_{\ell m}(\gamma_{a})R^{*}}{n^{2}},\quad n=1,2,\ldots,$
$\quad\ell=0,1,2,...n-1,\quad m=0,\pm 1,\pm 2,...\pm\ell,$ (22)
$R^{*}$ is the effective excitonic Rydberg energy
$R^{*}=\frac{\mu_{\parallel}e^{4}}{2(4\pi\epsilon_{0}\sqrt{\epsilon_{\parallel}\epsilon_{z}})^{2}\hbar^{2}},$ (23)
the anisotropy parameter $\gamma_{a}$ is defined as
$\gamma_{a}=\frac{\mu_{\parallel}\epsilon_{\parallel}}{\mu_{z}\epsilon_{z}},$
and $\eta_{\ell m}(\gamma_{a})$ is given by the following expression
@@
At room temperature, given the relatively low binding energy of excitons in Si (15 meV [26]), only the lowest excitonic state is relevant. Thus, one may assume $n=1,\ell=m=0$ and a single lifetime $T_{2n,n=1}=T_{2}$.
The solutions for $Y_{da,b\pm}^{(1)}$ determined above allow the calculation of the linear polarization
$P^{(1)}\left(\omega\right)=\int d^{3}{r}\left[Y_{da-}^{(1)}+Y_{da+}^{(1)*}\right]M^{*}({\bf r})$
$+\int{d}^{3}r\left[Y_{db-}^{(1)}+Y_{db+}^{(1)*}\right]M^{*}({\bf r})$
$=\frac{E_{0a}}{\hbar}\frac{2|c_{1}|^{2}\Omega_{1}}{\Omega^{2}_{1}-\left(\omega_{a}+{i}/T_{2}\right)^{2}}$
$+\frac{E_{0b}}{\hbar}\frac{2|c_{1}|^{2}\Omega_{1}}{\Omega^{2}_{1}-\left(\omega_{b}+{i}/T_{2}\right)^{2}}$
$=\epsilon_{0}\chi_{d}^{(1)}(\omega_{a})E_{0a}+\epsilon_{0}\chi_{d}^{(1)}(\omega_{b})E_{0b}.$ (24)
@@
The susceptibilities defined in equation (II.2) can be expressed in terms of the band parameters and, for energies below the gap, when spatial dispersion is neglected, we obtain
$\chi^{(1)}_{d}(\omega_{j})=\epsilon_{b}\frac{f_{1}\Delta_{LT}/R^{*}}{\left(E_{T1}-\hbar\omega_{j}-i{\mit\Gamma_{1}}\right)/R^{*}},$ (25)
@@
where ${\mit\Gamma}_{1}=\hbar/T_{2}$, and $E_{T}$ is the energy of the first exciton resonance. The anisotropy-dependent oscillator strength $f_{1}$ has the form
$f_{1}=\frac{\eta^{3}_{00}(1-\eta_{0}r_{0}/a^{*})}{(1+\eta_{00}r_{0}/a^{*})^{4}}.$ (26)
The above formula is obtained when the dipole density $M(\textbf{r})$ is assumed in the form
$M(\textbf{r})=M_{0}\,\frac{1}{\sqrt{4\pi}}\,\frac{1}{r\,r_{0}^{2}\gamma_{a}^{1/2}}e^{-r/r_{0}}Y_{00}(\theta,\phi),$ (27)
where $r_{0}=(2\mu E_{g}/\hbar^{2})^{-1/2}$ is the so-called coherence radius. The longitudinal transverse splitting of the ground state is
$\frac{\Delta_{LT}}{R^{*}}=2\frac{2\mu_{\parallel}}{\epsilon_{0}\epsilon_{b}\pi\,a^{*}\hbar^{2}}M_{0}^{2},$ (28)
@@
and is $f_{n}\Delta_{LT}$ for the excited states. Treating $\Delta_{LT}$ as a known quantity, the above relation determines the dipole matrix element $M_{0}$.
@@
To obtain the nonlinear response, the solutions for $Y_{da,b\pm}^{(1)}$ are inserted as a source term in the conduction band equation (9) and the valence band equation (10). Note that each of these equations depend on the electromagnetic field. If the irreversible terms are well defined, the equations (9,10) can be solved, and this second step of the iteration yields expressions for the density matrices $C$ and $D$. These density matrices can then, in turn, be used as a source term for equation (8), which can be solved to obtain expressions for $Y_{da,b\pm}^{(3)}$ in the final step of the iteration.
@@
Account for the presence of phonons, we obtain the nonlinear polarization for the region of discrete states. We separate the polarization related to the emission of a phonon (subscript “em”), and to the absorption of a phonon (subscript “abs”). The emission contribution has the form
$P^{(3)}_{d,em}(\omega_{a},\omega_{b})=P^{(3)}_{da,em}e^{-i\omega_{a}t}+P^{(3)}_{db,em}e^{-i\omega_{b}t},$ (29)
@@
$P^{(3)}_{da,em}=\epsilon_{0}\chi^{(3)}_{\hbox{\tiny d,self,em}}(\omega_{a},\omega_{a})|E(\omega_{a})|^{2}\,E(\omega_{a})$
$+\epsilon_{0}\chi^{(3)}_{\hbox{\tiny d,cross,em}}(\omega_{a},\omega_{b})|E(\omega_{b})|^{2}\,E(\omega_{a}),$
$P^{(3)}_{db,em}=\epsilon_{0}\chi^{(3)}_{\hbox{\tiny d, self,em}}(\omega_{b},\omega_{b})|E(\omega_{b})|^{2}\,E(\omega_{b})$
$+\epsilon_{0}\chi^{(3)}_{\hbox{\tiny d, cross,em}}(\omega_{b},\omega_{a})|E(\omega_{a})|^{2}\,E(\omega_{b}),$
@@
$\chi^{(3)}_{\hbox{ \tiny d,self,em}}(\omega_{j},\omega_{j})$
$=-(n_{ph}+1)\frac{2M_{0}^{2}}{\epsilon_{0}}\frac{1}{T_{2}}\left(T_{1}+\frac{i\hbar\delta(\omega-2\omega_{j})}{\hbar\omega+i\hbar/T_{1}}\right)$
×$\times\sum\limits_{\ell}\frac{c_{\ell}(A_{\ell}+B_{\ell})\hbar\Omega_{\ell,em}}{(\hbar\Omega_{\ell,em})^{2}-(\hbar\omega_{j}+i\hbar T_{2}^{-1})^{2}}$
×$\times\sum\limits_{n}\frac{c_{n}\varphi_{n}(0)}{(\hbar\Omega_{n,em}-\hbar\omega_{j})^{2}+(\hbar/T_{2})^{2}},$ (30)
$\chi^{(3)}_{\hbox{\tiny d,cross,em}}(\omega_{a},\omega_{b})=-(n_{ph}+1)\frac{2M_{0}^{2}}{\epsilon_{0}}\left(\frac{T_{1}}{T_{2}}\right)$
×$\times\sum\limits_{\ell}\frac{c_{\ell}(A_{\ell}+B_{\ell})\hbar\Omega_{\ell,em}}{(\hbar\Omega_{\ell,em})^{2}-(\hbar\omega_{a}+i\hbar T_{2}^{-1})^{2}}$
×$\times\sum\limits_{n}\frac{c_{n}\varphi_{n}(0)}{(\hbar\Omega_{n,em}-\hbar\omega_{b})^{2}+(\hbar/T_{2})^{2}},$ (31)
and the additional cross term $\chi^{(3)}_{\hbox{\tiny d,cross,em}}(\omega_{b},\omega_{a})$ is obtained by permuting the frequencies $\omega_{a}$ and $\omega_{b}$ in equation (31). The number of available phonons $n_{ph}$ is estimated by the formula
$n_{ph}=\frac{1}{|\exp\frac{\hbar\omega_{ph}}{k_{B}{\mathcal{T}}}-1|},$ (32)
with $k_{B}$ is the Boltzmann constant, and $\mathcal{T}$ the temperature. $\hbar\Omega_{n,em,abs}$ are the exciton resonance energies that include the phonon energies $\hbar\omega_{ph}$,
$\hbar\Omega_{n,em,abs}=E_{g}+E_{n}\pm\hbar\omega_{ph},$ (33)
where “+” stands for phonon emission and “-” for phonon absorption. As discussed in [27, 28], the phonon density of states contains two local maxima at 20 meV and 60 meV, respectively, with a weighted average of approximately 40 meV. This simplified approach of taking an estimate of the average phonon energy has been used in previous calculations, providing proper fits to experimental data.
The coefficients $A_{\ell},B_{\ell}$ in equations (30,31) are related to the irreversible terms in the equations (9,10), and have the form
$A_{\ell}=\int d^{3}r\;\varphi_{\ell}(\textbf{r})\;e^{-r^{2}/2\lambda_{th,e}^{2}},\quad\ell=1,2,\ldots,$
$B_{\ell}=\int d^{3}r\,\varphi_{\ell}(\textbf{r})\;e^{-r^{2}/2\lambda_{th,h}^{2}},$ (34)
where $\lambda_{th,e},\lambda_{th,h}$ are the so-called thermal lengths for electrons and holes, respectively,
$\lambda_{th,e}=\left(\frac{\hbar^{2}}{m_{e}k_{B}{\mathcal{T}}}\right)^{1/2},\;\;\lambda_{th,h}=\left(\frac{\hbar^{2}}{m_{h}k_{B}{\mathcal{T}}}\right)^{1/2}.$
The above expressions are valid when
$\hbar\omega<E_{g}+\hbar\omega_{ph}.$
For $\hbar\omega>E_{g}+\hbar\omega_{ph}$ one should replace $n_{ph}+1$ by
@@
Analogous expressions can be obtained for the susceptibilities related to the phonon absorption
$\chi^{(3)}_{\hbox{\tiny d, self,abs}}(\omega_{j},\omega_{j})=-n_{ph}\frac{2M_{0}^{2}}{\epsilon_{0}}\frac{1}{T_{2}}\left(T_{1}+\frac{i\hbar\delta(\omega-2\omega_{j})}{\hbar\omega+i\hbar/T_{1}}\right)$
×$\times\sum\limits_{\ell}\frac{c_{\ell}(A_{\ell}+B_{\ell})\hbar\Omega_{\ell,abs}}{(\hbar\Omega_{\ell,abs})^{2}-(\hbar\omega_{j}+i\hbar T_{2}^{-1})^{2}}$
×$\times\sum\limits_{n}\frac{c_{n}\varphi_{n}(0)}{(\hbar\Omega_{n,abs}-\hbar\omega_{j})^{2}+(\hbar/T_{2})^{2}},$ (36)
$\chi^{(3)}_{\hbox{\tiny d, cross,abs}}(\omega_{a},\omega_{b})=-n_{ph}\frac{2M_{0}^{2}}{\epsilon_{0}}\left(\frac{T_{1}}{T_{2}}\right)$
×$\times\sum\limits_{\ell}\frac{c_{\ell}(A_{\ell}+B_{\ell})\hbar\Omega_{\ell,abs}}{(\hbar\Omega_{\ell,abs})^{2}-(\hbar\omega_{a}+i\hbar T_{2}^{-1})^{2}}$
×$\times\sum\limits_{n}\frac{c_{n}\varphi_{n}(0)}{(\hbar\Omega_{n,abs}-\hbar\omega_{b})^{2}+(\hbar/T_{2})^{2}},$ (37)
plus the additional cross term $\chi^{(3)}_{\hbox{\tiny d, cross,abs}}(\omega_{b},\omega_{a})$ obtained by permuting the input frequencies in (37). The above expressions are valid when

### v2

@@ II.3 Discrete states - linear susceptibility
$\left(\frac{\partial Y^{(1)}}{\partial t}\right)_{{\rm irrev}}=-\frac{1}{T_{2}}Y^{(1)}.$ (46)
In the discussion of nonlinear effects we also take into account the non-resonant parts of the amplitude $Y$. The excitonic density $Y$ will consists of two parts, $Y_{a},Y_{b}$, as defined by the angular frequencies $\omega_{a}$ and $\omega_{b}$. In addition, due to the valence band structure of the semiconducting material (Si) we must consider heavy-hole (H) and light hole (L) excitons. Considering optical transitions between the (H,L) valence bands and the conduction band, with the mentioned inclusion of both the resonant and anti-resonant parts, Eq. (12) generates eight equations: a pair for amplitude $Y_{aH}$: $Y_{aH-}^{(1)}\,\propto\exp(-{i}\omega_{a}t)$ and for $Y^{(1)}_{aH+}\,\propto\exp({i}\omega_{a}t)$
${i}\hbar\left(i\omega_{a}+\frac{1}{T_{2}}\right)Y^{(1)}_{aH+}-H_{eh}Y^{(1)}_{aH+}$
$=-{\bf M}_{H}{\bf E}_{a}^{*}({\bf R},t),$ (47)
${i}\hbar\left(-{i}\omega_{a}+\frac{1}{T_{2}}\right)Y^{(1)}_{aH-}-H_{eh}Y^{(1)}_{aH-}$ (48)
$=-{\bf M}_{H}{\bf E}_{a}({\bf R},t),$
with similar equations for $Y_{bH\pm}^{(1)}$, where $\textbf{M}_{H}$ is the transition dipole density. Analogous equations hold for the amplitudes $Y_{a,bL\pm}^{(1)}$, with the appropriate transition dipole density $\textbf{M}_{L}$. In the following, we consider only one component of the vectors ${\bf E}$, ${\bf P}$ and ${\bf M}$, and focus our attention on the heavy hole exciton transition.
@@
$Y^{(1)}_{daH-}=E_{a}({\bf R},t)\sum_{n\ell m}\frac{c_{n\ell mH}\varphi_{n}({\bf r})}{\hbar(\Omega_{n\ell mH}-\omega_{a}-{i}/T_{2})},$
$Y^{(1)}_{daH+}=E_{a}^{*}({\bf R},t)\sum_{n\ell m}\frac{c_{n\ell m}\varphi_{n\ell mH}({\bf r})}{\hbar(\Omega_{n\ell mH}+\omega_{a}-{i}/T_{2})},$ (49)
and similar expressions for $Y_{dbH\pm}^{(1)}$. The subscript “d” indicates the case of discrete excitonic states. The expansion coefficients are defined as follows
$c_{n\ell mH}=\int{d}^{3}rM_{H}({\bf r})\varphi_{n\ell mH}({\bf r}),$
$\hbar\Omega_{n\ell mH}=\hbar\Omega_{n\ell mH}=E_{g}+E_{n\ell mH}(\gamma_{aH}),$ (50)
$\varphi_{n\ell mH}=R_{n\ell H}(r)Y_{\ell m}(\theta,\phi),$
$R_{n\ell H}$ are the hydrogen radial functions of the anisotropic Schrödinger equation [39],
$r=\sqrt{x^{2}+y^{2}+\frac{z^{2}}{\gamma_{aH}}},$ (51)
$Y_{\ell m}(\theta,\phi)$ are the spherical harmonics, and $E_{n\ell mH}$ the corresponding eigenvalues
$\phantom{nuc}E_{n\ell mH}=-\frac{\eta^{2}_{\ell mH}(\gamma_{aH})R_{H}^{*}}{n^{2}},\quad n=1,2,\ldots,$
$\quad\ell=0,1,2,...n-1,\quad m=0,1,2,...\ell,$ (52)
In the above, $R_{H}^{*}$ is the effective excitonic Rydberg energy for the heavy hole exciton
$R_{H}^{*}=\frac{\mu_{\parallel H}e^{4}}{2(4\pi\epsilon_{0}\sqrt{\epsilon_{\parallel}\epsilon_{z}})^{2}\hbar^{2}},$ (53)
and the anisotropy parameter $\gamma_{aH}$ is defined as
$\gamma_{aH}=\frac{\mu_{\parallel H}\epsilon_{\parallel}}{\mu_{zH}\epsilon_{z}},$
where $\mu_{\parallel H}$ and $\mu_{zH}$ are the heavy hole exciton reduced masses in the $x-y$ plane and in the $z$-direction, respectively,
$\frac{1}{\mu_{\parallel H}}=\frac{1}{m_{e\parallel}}+\frac{1}{m_{h\parallel H}},$
$\frac{1}{\mu_{zH}}=\frac{1}{m_{ez}}+\frac{1}{m_{hzH}}.$ (54)
The anisotropic electron and heavy hole masses (in-plane and in the $z$-direction) are listed in Table 1, $\epsilon_{0}$ is the vacuum dielectric constant, and $\epsilon_{\parallel},\epsilon_{z}$ are relative dielectric tensor elements. The quantity $\eta_{\ell mH}(\gamma_{aH})$ is given by the following expression
@@
In what follows we consider only the excitonic $s$ states, with $\ell=m=0$, and denote
$\varphi_{n00H}=\varphi_{nH},\quad c_{n00H}=c_{nH},$
$E_{n00H}=E_{nH},\quad\hbar\Omega_{n00H}=\hbar\Omega_{nH}+E_{n}(\gamma_{aH}).$ (56)
@@
It should be noted that at room temperature, accounting for the relatively low binding energy of excitons in Si (15 meV as in Ref. [40]), only the lowest excitonic state is relevant, so we set $n=1$.
The solutions for $Y_{da,b\pm}^{(1)}$ determined above permit the calculation of the linear polarization
$P_{H}^{(1)}\left(\omega\right)=\int d^{3}{r}\left[Y_{daH-}^{(1)}+Y_{daH+}^{(1)*}\right]M_{H}^{*}({\bf r})$
$+\int{d}^{3}r\left[Y_{dbH-}^{(1)}+Y_{dbH+}^{(1)*}\right]M_{H}^{*}({\bf r})$
$=\frac{E_{0a}}{\hbar}\frac{2|c_{1H}|^{2}\Omega_{1H}}{\Omega^{2}_{1H}-\left(\omega_{a}+{i}/T_{2}\right)^{2}}$
$+\frac{E_{0b}}{\hbar}\frac{2|c_{1H}|^{2}\Omega_{1H}}{\Omega^{2}_{1H}-\left(\omega_{b}+{i}/T_{2}\right)^{2}}$
$=\epsilon_{0}\chi_{dH}^{(1)}(\omega_{a})E_{0a}+\epsilon_{0}\chi_{dH}^{(1)}(\omega_{b})E_{0b}.$ (57)
@@
The susceptibilities defined in equation (II.3) can be expressed in terms of the band parameters and, for energies below the gap, when spatial dispersion is neglected, we obtain
$\chi^{(1)}_{dH}(\omega_{j})=\epsilon_{b}\frac{f_{1H}\Delta_{LTH}/R_{H}^{*}}{\left(E_{T1H}-\hbar\omega_{j}-i\hbar/T_{2}\right)/R_{H}^{*}},$ (58)
where $j=a,b$, and $E_{T1H}$ is the energy of the first heavy hole exciton resonance. For the dipole density $M_{H}(\textbf{r})$ described by the formula
$M_{H}(\textbf{r})$ (59)
$=M_{0H}\,\frac{1}{\sqrt{4\pi}}\,\frac{1}{r\,r_{0H}^{2}\gamma_{aH}^{1/2}}e^{-r/r_{0H}}Y_{00}(\theta,\phi),$
@@
where $r_{0H}=(2\mu_{\parallel H}E_{g}/\hbar^{2})^{-1/2}$ is the so-called coherence radius, the anisotropy-dependent oscillator strength $f_{1H}$ has the form
$f_{1H}=\frac{\eta^{3}_{00H}(1-\eta_{00H}r_{0H}/a_{H}^{*})}{(1+\eta_{00H}r_{0H}/a_{H}^{*})^{4}},$ (60)
where $a_{H}^{*}$ is the heavy hole effective excitonic Bohr radius
$a_{H}^{*}=\frac{4\pi\hbar^{2}\epsilon_{0}\sqrt{\epsilon_{\parallel}\epsilon_{z}}}{\mu_{\parallel H}e^{2}}.$ (61)
The longitudinal transverse splitting of the ground state is [29]
$\frac{\Delta_{LTH}}{R^{*}}=2\frac{2\mu_{\parallel H}}{\epsilon_{0}\epsilon_{b}\pi\,a_{H}^{*}\hbar^{2}}M_{0H}^{2},$ (62)
@@
where the above mentioned bulk dielectric constant is given by $\epsilon_{b}=\sqrt{\epsilon_{z}\epsilon_{\parallel}}$, and is $f_{n}\Delta_{LTH}$ for the excited states ($n>1$). Treating $\Delta_{LTH}$ as a known quantity, the above relation allows computation of the dipole matrix element $M_{0H}$.
II.4 Discrete states - Nonlinear susceptibility
@@
To obtain the nonlinear response, the solutions for $Y_{daH,b\pm}^{(1)}$ are inserted as a source term in the conduction band equation (24) and the valence band equation (25). Note that each of these equations depends on the electromagnetic field. If the irreversible terms are well defined, the equations (24,25) can be solved, and this second step of the iteration yields expressions for the density matrices $C_{H}$ and $D_{H}$. We use for the irreversible terms a linear relaxation time approximation
$\left(\frac{\partial C}{\partial t}\right)_{\rm irrev}$ (63)
$=-\frac{1}{\tau}\left[C({\bf X},{\bf r},t)-f_{0e}({\bf r})C({\bf X},{\bf r}=0,t)\right]-\frac{C(0)}{T_{1}},$
where ${\bf X}=({\bf r}_{1}+{\bf r}_{2})/2$, $\tau$ is the carrier relaxation time, and $f_{0e},f_{0h}$ are normalized Boltzmann distributions for electrons and holes, respectively,
$f_{0e}({\bf r})=\int d^{3}{q}f_{0e}({\bf q})e^{-i{\bf qr}}$ (64)
$=\exp\left(-\frac{m_{e\parallel}k_{B}{\mathcal{T}}}{2\hbar^{2}}\rho^{2}-\frac{m_{ez}\hbox{kT}}{2\hbar^{2}}z^{2}\right),$
The same type of expression holds for the holes. The diagonal elements of the matrices are related to charge densities (12, 13), which are conserved quantities. Therefore, we assume that they relax to an equilibrium normalized to the actual number of carriers. The relaxation time $T_{1}$ stands for the interband recombination time. [41] Using the irreversible terms (63) in the intraband equations (24, 25), and looking for stationary solutions, we obtain the matrices $C,D$ in the form
$C_{H}({\bf r})=-\frac{i}{\hbar}\left[\tau J_{CH}({\bf r})-\tau J_{CH}(0)+T_{1}f_{0e}({\bf r})J_{CH}(0)\right],$
$D_{H}({\bf r})=-\frac{\rm i}{\hbar}\biggl{[}\tau J_{HH}({\bf r})$
$-\tau J_{HH}(0)+T_{1H}f_{0hH}({\bf r})J_{HH}(0)\biggr{]},$
where
$J_{CH}=\frac{2{i}M_{0}|{E}_{0a}|^{2}}{\hbar}\biggl{[}\hbox{Im}\,g_{H}(-\omega_{a},{\bf r})+\,\hbox{Im}\,g_{H}(\omega_{a},{\bf r})\biggr{]}$
$+\frac{2{i}M_{0}|{E}_{0b}|^{2}}{\hbar}\biggl{[}\hbox{Im}\,g_{H}(-\omega_{b},{\bf r})+\,\hbox{Im}\,g_{H}(\omega_{b},{\bf r})\biggr{]},$ (65)
and
$g_{H}(\pm\omega_{j},{\bf r})=\sum_{n}\frac{c_{nH}\varphi_{nH}({\bf r})}{\Omega_{nH}\mp\omega_{j}-{i}/T_{2}},$ (66)
with $J_{CH}=J_{DH}$. These density matrices can then, in turn, be used as a source term for equation (23), which can be solved to obtain expressions for $Y_{da,bH\pm}^{(3)}$ in the final step of the iteration.
The equations for the third-order coherent amplitudes $Y^{(3)}_{jH\pm}$ now take on the form
${i}\hbar\left(i\omega_{a}+\frac{1}{T_{2}}\right)Y^{(3)}_{aH+}-H_{eh}Y^{(3)}_{aH+}$
$=M_{0H}({E}^{*}C_{H}+{E}^{*}D_{H})={E}^{*}({\bf R},t)\tilde{J}_{H},$ (67)
${i}\hbar\left(-{i}\omega_{a}+\frac{1}{T_{2}}\right)Y^{(3)}_{aH-}-H_{eh}Y^{(3)}_{aH-}$ (68)
$=M_{0H}({E}^{*}C_{H}+{E}^{*}D_{H})={E}^{*}({\bf R},t)\tilde{J}_{H},$
with similar equations for $Y_{bH\pm}^{(1)}$, where
$\tilde{J}_{H}=-\frac{i}{\hbar}M_{0H}\left[T_{1}J_{CH}(0)f_{0eH}({\bf r})\right.$
$\left.+T_{1}J_{VH}(0)f_{0hH}(\textbf{r})\right].$ (69)
Once $Y^{(3)}_{aH\pm}$ and $Y^{(3)}_{bH\pm}$ are know, the third-order polarization can be determined as
$P_{H}^{(3)}\left(\omega\right)=\int d^{3}{r}\left[Y_{aH-}^{(3)}+Y_{aH+}^{(3)*}\right]M_{H}^{*}({\bf r})$
$+\int{d}^{3}r\left[Y_{bH-}^{(3)}+Y_{bH+}^{(3)*}\right]M^{*}_{H}({\bf r}).$ (70)
We consider nonlinear polarizations at the same frequency $\omega$ (or frequencies in the case 2PA) as that of an incident field, which means that we consider the susceptibilities $\chi^{(1)}$ and $\chi^{(3)}$ related to the excitonic amplitudes $Y^{(1)}$ and $Y^{(3)}$, respectively.
@@
Similar to the approach presented in Section II.1, to account for the presence of phonons, we separate the polarization related to the emission of a phonon (subscript “em”) and the absorption of a phonon (subscript “abs”). The emission contribution has the form
$P^{(3)}_{dH,em}(\omega_{a},\omega_{b})$ (71)
$=P^{(3)}_{daH,em}e^{-i\omega_{a}t}+P^{(3)}_{dbH,em}e^{-i\omega_{b}t},$
@@
$P^{(3)}_{daH,em}=\epsilon_{0}\chi^{(3)}_{\hbox{\tiny d H,self,em}}(\omega_{a},\omega_{a})|E(\omega_{a})|^{2}\,E(\omega_{a})$
$+\epsilon_{0}\chi^{(3)}_{\hbox{\tiny d H,cross,em}}(\omega_{a},\omega_{b})|E(\omega_{b})|^{2}\,E(\omega_{a}),$
$P^{(3)}_{dbH,em}=\epsilon_{0}\chi^{(3)}_{\hbox{\tiny d H, self,em}}(\omega_{b},\omega_{b})|E(\omega_{b})|^{2}\,E(\omega_{b})$
$+\epsilon_{0}\chi^{(3)}_{\hbox{\tiny d H, cross,em}}(\omega_{b},\omega_{a})|E(\omega_{a})|^{2}\,E(\omega_{b}),$
@@
$\chi^{(3)}_{\hbox{ \tiny d H,self,em}}(\omega_{j},\omega_{j})$
$=-(n_{ph}+1)\frac{2M_{0H}^{2}}{\epsilon_{0}}\frac{1}{T_{2}}\left(T_{1}+\frac{i\hbar\delta(\omega-2\omega_{j})}{\hbar\omega+i\hbar/T_{1}}\right)$
×$\times\sum\limits_{\ell}\frac{c_{\ell H}(A_{\ell H}+B_{\ell H})\hbar\Omega_{\ell H,em}}{(\hbar\Omega_{\ell H,em})^{2}-(\hbar\omega_{j}+i\hbar T_{2}^{-1})^{2}}$ (72)
×$\times\sum\limits_{n}\frac{c_{nH}\varphi_{nH}(0)}{(\hbar\Omega_{nH,em}-\hbar\omega_{j})^{2}+(\hbar/T_{2})^{2}},$
$\chi^{(3)}_{\hbox{\tiny d H,cross,em}}(\omega_{a},\omega_{b})=-(n_{ph}+1)\frac{2M_{0H}^{2}}{\epsilon_{0}}\left(\frac{T_{1}}{T_{2}}\right)$
×$\times\sum\limits_{\ell}\frac{c_{\ell H}(A_{\ell H}+B_{\ell H})\hbar\Omega_{\ell H,em}}{(\hbar\Omega_{\ell H,em})^{2}-(\hbar\omega_{a}+i\hbar T_{2}^{-1})^{2}}$ (73)
×$\times\sum\limits_{n}\frac{c_{nH}\varphi_{nH}(0)}{(\hbar\Omega_{nH,em}-\hbar\omega_{b})^{2}+(\hbar/T_{2})^{2}}.$
The additional cross term $\chi^{(3)}_{\hbox{\tiny d H,cross,em}}(\omega_{b},\omega_{a})$ is obtained by permuting the frequencies $\omega_{a}$ and $\omega_{b}$ in equation (71). The expressions $\hbar\Omega_{nH,em,abs}$ are the exciton resonance energies that include the phonon energies $\hbar\omega_{ph}$
$\hbar\Omega_{nH,em,abs}=E_{g}+E_{nH}\pm\hbar\omega_{ph},$ (74)
where “+” stands for phonon emission and “-” for phonon absorption. The coefficients $A_{\ell H},B_{\ell H}$ in equations (II.4) appearing in the expressions for $Y_{da,bH\pm}^{(3)}$ have the form
$A_{\ell H}=\int d^{3}r\;\varphi_{\ell H}(\textbf{r})\;f_{0e}({\bf r})$
$=\int d^{3}r\;\varphi_{\ell H}(\textbf{r})\exp\left(-\frac{\rho^{2}}{2\lambda_{th,e\parallel}^{2}}-\frac{z^{2}}{2\lambda_{th,ez}^{2}}\right),$
$B_{\ell H}=\int d^{3}r\,\varphi_{\ell H}(\textbf{r})\;f_{0hH}({\bf r})$
$\int d^{3}r\;\varphi_{\ell H}(\textbf{r})\exp\left(-\frac{\rho^{2}}{2\lambda_{th,h\parallel H}^{2}}-\frac{z^{2}}{2\lambda_{th,hzH}^{2}}\right),$
where $\lambda_{th,e},\lambda_{th,hH}$ are the so-called thermal lengths for electrons and holes, respectively,
$\lambda_{th,e}=\left(\frac{\hbar^{2}}{m_{e}k_{B}{\mathcal{T}}}\right)^{1/2},\;\;\lambda_{th,hH}=\left(\frac{\hbar^{2}}{m_{h}k_{B}{\mathcal{T}}}\right)^{1/2},$ (76)
determined for the appropriate masses ($\parallel$ or $z$). The above expressions are valid when $\hbar\omega<E_{g}+\hbar\omega_{ph}$. For $\hbar\omega>E_{g}+\hbar\omega_{ph}$ one should replace $n_{ph}+1$ by
@@
Analogous expressions can be obtained for the susceptibilities that include phonon absorption
$\chi^{(3)}_{\hbox{\tiny dH, self,abs}}(\omega_{j},\omega_{j})=-n_{ph}\frac{2M_{0}^{2}}{\epsilon_{0}}\frac{1}{T_{2}}\left(T_{1}+\frac{i\hbar\delta(\omega-2\omega_{j})}{\hbar\omega+i\hbar/T_{1}}\right)$
×$\times\sum\limits_{\ell}\frac{c_{\ell}(A_{\ell H}+B_{\ell H})\hbar\Omega_{\ell H,abs}}{(\hbar\Omega_{\ell H,abs})^{2}-(\hbar\omega_{j}+i\hbar T_{2}^{-1})^{2}}$
×$\times\sum\limits_{n}\frac{c_{nH}\varphi_{nH}(0)}{(\hbar\Omega_{nH,abs}-\hbar\omega_{j})^{2}+(\hbar/T_{2})^{2}},$ (78)
$\chi^{(3)}_{\hbox{\tiny dH, cross,abs}}(\omega_{a},\omega_{b})=-n_{ph}\frac{2M_{0H}^{2}}{\epsilon_{0}}\left(\frac{T_{1}}{T_{2}}\right)$
×$\times\sum\limits_{\ell}\frac{c_{\ell H}(A_{\ell H}+B_{\ell})\hbar\Omega_{\ell H,abs}}{(\hbar\Omega_{\ell H,abs})^{2}-(\hbar\omega_{a}+i\hbar T_{2}^{-1})^{2}}$
×$\times\sum\limits_{n}\frac{c_{nH}\varphi_{nH}(0)}{(\hbar\Omega_{nH,abs}-\hbar\omega_{b})^{2}+(\hbar/T_{2})^{2}},$ (79)
plus the additional cross term $\chi^{(3)}_{\hbox{\tiny dH, cross,abs}}(\omega_{b},\omega_{a})$ obtained by permuting the input frequencies in (79). The above expressions are valid when